### v1

# Benchmarking MedMNIST dataset on real quantum hardware

Abstract
Quantum machine learning (QML) has emerged as a promising domain to leverage the computational capabilities of quantum systems to solve complex classification tasks. In this work, we present first comprehensive QML study by benchmarking the MedMNIST-a diverse collection of medical imaging datasets on a 127-qubit real IBM quantum hardware, to evaluate the feasibility and performance of quantum models (without any classical neural networks) in practical applications. This study explore recent advancements in quantum computing such as device-aware quantum circuits, error suppression and mitigation for medical image classification. Our methodology comprised of three stages: preprocessing, generation of noise-resilient and hardware-efficient quantum circuits, optimizing/training of quantum circuits on classical hardware, and inference on real IBM quantum hardware. Firstly, we process all input images in the preprocessing stage to reduce the spatial dimension due to the quantum hardware limitations. We generate hardware-efficient quantum circuits using backend properties expressible to learn complex patterns for medical image classification. After classical optimization of QML models, we perform the inference on real quantum hardware. We also incorporates advanced error suppression and mitigation techniques in our QML workflow including dynamical decoupling (DD), gate twirling, and matrix-free measurement mitigation (M3) to mitigate the effects of noise and improve classification performance. The experimental results showcase the potential of quantum computing for medical imaging and establishes a benchmark for future advancements in QML applied to healthcare.
Keywords Quantum machine learning  $\cdot$ Medical image classification  $\cdot$ Device-aware quantum circuit  $\cdot$ Error mitigation
1 Introduction
Biomedical image analysis plays a critical role in advancing human healthcare. Recently, machine learning (ML) has demonstrated remarkable potential in this domain, excelling in tasks such as medical image classification [1, 2, 3], segmentation [4, 5], and reconstruction [6, 7]. Simultaneously, QML an emerging paradigm of ML implemented on quantum computers, is anticipated to offer quantum advantages over classical ML. These advantages include enhanced speed-ups, improved model expressivity, efficient exploration of high-dimensional feature spaces, and the ability to address certain classically intractable problems [8, 9, 10, 11, 12]. QML capitalizes on the unique properties of quantum mechanics, such as superposition, entanglement, and interference, to unravel complex relationships between inputs and labels. In the current noisy intermediate-scale quantum (NISQ) era, QML faces notable challenges. Limited access to large-scale quantum systems constrains the ability to encode high-dimensional data into quantum states. Hardware noise and decoherence further impact accuracy and restrict circuit depth, thereby limiting the expressivity of QML models. Rapid advancements in quantum hardware, including improved quantum error correction techniques [13], are expected to overcome these limitations and enable the practical deployment of QML for real-world applications in the near future.
Despite these obstacles, QML research has witnessed significant growth [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Generally, QML approaches are based on parameterized quantum circuits (PQCs) [29] consisting of fixed gates (such as SWAP and CNOT) and single-qubit (such as Pauli rotation) gates with trainable parameters. These PQCs comprised of three stages: data encoding, a variational ansatz circuit with trainable parameters, and measurement to classical bits. The data encoding stage encode the classical input data to quantum states. The parameters of the ansatz circuit are optimized using classical gradient descent algorithms, typically relying on the parameter-shift rule [30], enabling the circuit to solve machine learning tasks.
Over the last few years, researchers have explored QML algorithms for medical imaging in various studies including [26, 31, 32, 33, 34, 35, 36, 37, 38, 39] to harness potential quantum advantages over classical machine learning approaches. These studies address a variety of medical imaging modalities, including MRI, CT, X-ray, OCT, and ultrasound, targeting tasks such as clinical data classification and detection. The QML algorithms investigated in these works include quantum support vector machines (SVM) [34], hybrid quantum-classical neural networks [26, 35, 36, 38, 39], quantum transfer learning [33, 37], and variational quantum circuits [31, 32]. To the best of our knowledge, few studies such as [26, 33, 34, 37, 39] have used real quantum hardware for evaluation purposes but at a limited scale. Cherrat et al. proposed a quantum vision transformer (QViT) in [26] based on shallow PQCs capable of performing complex image classification tasks. This QViT introduces trainable quantum orthogonal layers adaptable to varying levels of quantum hardware connectivity. The authors have evaluated their QViT on open source MedMNIST (a collection of 12 biomedical imaging datasets) [1] using a simulator and real hardware results are provided for only one dataset (RetinaMNIST) out of 12. In [33], Acar and Yilmaz designed a hybrid classical-quantum transfer learning based QML model for COVID-19 detection using CT images. The authors claimed to achieve a classification accuracy of 94-100% on quantum devices with limited qubit capacity compared to 90% on classical systems using a small dataset. This highlights the potential of quantum computers for efficient and accurate medical image classification, particularly with limited data. Moradi et al. [34] designed two QML methods i.e., quantum distance classifier and quantum-kernel SVM for clinical data classification on quantum hardware. The findings observed in this study also highlight the promise of QML in clinical data analysis, especially for tasks constrained by dataset size and feature complexity. Azevedo et al. [37] employed hybrid quantum-classical neural network for mammogram classification into malignant and benign categories. The authors compared the results obtained from quantum device and simulator, highlighting the potential of QML in medical diagnostics while emphasizing the need for further validation. Similarly, the study conducted in [39] by Landman et al. introduces two quantum neural networks: quantum-assisted neural networks, which leverage quantum computer to estimate inner products during the training and inference of classical neural networks, and quantum orthogonal neural networks, built on quantum pyramidal circuits for orthogonal matrix multiplication. This study also considered only two datasets (PneumoniaMNIST and RetinaMnist) out of 12 datasets from MedMNIST for evaluation purposes on real hardware. Moreover, all of these studies showing evaluation on quantum hardwares are small-scale devices with 5, 7, 15, 16, and 27 qubits. It is also worth noting that most of these QML studies are based on hybrid quantum-classical neural networks which combines quantum circuits with classical neural networks, leaving the potential of purely quantum approaches underexplored. Also, the individual contribution of both quantum and classical parts is unclear. While some studies claim potential quantum advantages, these are often observed only in restricted settings (e.g., small datasets or low-qubit systems) with limited evidence of their generalization across different image classification domains or modalities. Demonstrating clear, reproducible quantum advantages over classical machine learning models remains a significant challenge.
To address above mentioned limitations, we conducted a comprehensive QML study for medical image classification on a 127 qubits IBM quantum hardware. This research study aims to explore the different aspects of QML for image classification including efficient encoding of high-dimensional image data to quantum states, generation of device-aware quantum circuits having optimal expressivity for different image datasets, advantages of error suppression and mitigation techniques. By exclusively employing quantum methods without classical neural networks, our work addresses critical gaps in the current QML literature, offering insights into the practical applicability and performance of quantum-only models for medical image classification tasks. This effort not only highlights the promise of QML in addressing real-world problems but also sets a foundational benchmark for future explorations in this domain. The main contributions of our research study are as follows:
  • We present the first comprehensive QML study by benchmarking the MedMNIST dataset on real IBM Cleveland quantum hardware.
  • This research leverages and evaluates the recent quantum computing capabilities such as generation of device-aware quantum circuits and error suppression/mitigation techniques for medical image classification.
  • We also conducted an ablation study on different error suppression and mitigation techniques to provide an intuitive understanding.
  • The experimental results provide foundational insights into the performance and scalability of quantum models for medical image classification, establishing a benchmark for future QML research for healthcare applications.
In the remaining sections of this research study, we firstly discussed our QML methodology in detail in Section 2. Then, we provide the experiment results with discussion in Section 3. Lastly, concluding remarks are provided in Section 4.
2 Methodology
In this section, we discuss our QML methodology for medical image classification, as illustrated in Fig. 1. Our approach is structured into four key stages: data preprocessing, generation of device-aware quantum circuits, training of the quantum circuits on classical hardware using training dataset, and inference of the trained quantum model on real IBM quantum hardware using a testing dataset. It is important to emphasize that, due to the limitations of current NISQ quantum devices, training quantum models with large datasets directly on hardware remains infeasible. To address this challenge, our QML workflow adopts a hybrid strategy, where the training phase is executed on classical hardware, leveraging their computational power, while the inference stage is performed on real quantum hardware to validate the model’s performance. The following subsections provide a detailed explanation of each stage of our QML methodology.
Figure 1: QML framework for medical image classification.
2.1 Data preprocessing
The preprocessing stage focuses on reducing the feature dimensionality of input images, a necessary step in our QML workflow due to the constraints of current quantum hardware. Specifically, we employ angle encoding to map classical data to quantum states. The angle encoding method is limited by the hardware’s inability to handle high-dimensional feature spaces in the current NISQ era. Therefore, it is very important to reduce the spatial dimension of the images effectively. We used average pooling to downscale the input image dimension from 224$\times$224 to 7$\times$7 and 8$\times$8 for the considered MedMNIST datasets, thereby providing 49 and 64 features, respectively. This indicates that performing angle encoding for 49 and 64 features requires the incorporation of 49 and 64 Pauli rotation gates, respectively, within the quantum circuit. The quantum circuit representational capacity should be adjusted accordingly to effectively capture the patterns in the input features. Post-feature reduction, the pixel values are normalized to the range $[0,\pi]$, aligning with the requirements of angle encoding for efficient quantum state preparation. It is important to note that this preprocessing stage introduces information loss compared to classical machine learning models, which can process entire images as input without dimensionality reduction. In QML workflow, this limitation arises due to the constraints of current quantum hardware, which restrict the number of features that can be encoded into quantum states. Consequently, while classical models can leverage the full spatial resolution of input images, quantum models operating with a reduced feature set, potentially impacting the overall representational capacity. This trade-off underscores the need for effective feature reduction techniques that preserve essential information while ensuring compatibility with the limitations of quantum devices.
2.2 Efficient and noise-resilient quantum circuits for QML
The second crucial step of our QML framework involves designing a noise-robust and efficient quantum circuit that is expressively powerful enough to learn meaningful patterns from the input data for medical imaging classification task. In the current NISQ era, developing device-aware quantum circuits tailored to the target backend properties, such as qubit connectivity and gate fidelities, has emerged as a key strategy to mitigate the impact of hardware noise. However, designing such circuits remains a significant challenge due to the lack of well-defined guidelines for quantum circuit architecture in QML workflows. Additionally, the optimal design of a quantum circuit is highly dataset-dependent, as varying datasets require different levels of expressivity to effectively capture their underlying patterns.
To address these challenges, we incorporated the Élivágar framework [40] into our QML workflow to generate hardware-aware quantum circuits tailored to our medical image classification task. This framework allowed us to create noise-resilient and computationally efficient quantum circuits that align with the physical constraints of the target quantum hardware while maximizing model performance. The quantum circuits in our workflow comprise three key components: (1) data encoding gates, (2) a variational or ansatz circuit, and (3) measurement operations that map quantum states to classical bits. The data encoding layer leverages parameterized rotation gates to embed classical input data into quantum states. The ansatz, or variational circuit, contains trainable Pauli rotation gates interspersed with CNOT gates to ensure qubit entanglement. The parameters of these gates ($\theta$) are optimized during training to minimize the loss function. Input data from the preprocessing stage is encoded using a number of rotation gates equal to the feature count of each input sample. An essential step in the Élivágar framework is computing the Clifford Noise Resilience (CNR) to assess the robustness of candidate circuits against hardware-induced errors. While this process is typically conducted on quantum device, it is time-intensive due to the need to run thousands of circuits, often facing significant queuing delays. To streamline the process, we performed CNR calculations using a simulator, enabling efficient evaluation of circuit noise resilience while saving considerable time.
To ensure adaptability and effectiveness, we designed multiple quantum circuits with varying levels of expressivity by systematically adjusting the number of trainable parameters and the depth of the ansatz. These circuits were carefully tailored to meet the unique requirements of the MedMNIST datasets. Specifically, we generated quantum circuits with different configurations, incorporating varying numbers of trainable parameters—such as 60, 80, 100, and 120 for all the MedMNIST datasets under consideration. This approach allowed us to explore the impact of circuit expressivity on the performance of our quantum models and ensured optimal alignment with the specific characteristics of each dataset. Importantly, to thoroughly investigate the potential of quantum circuits in medical imaging classification, our QML framework exclusively employs quantum models without incorporating any classical neural network components. This approach highlights the intrinsic capabilities of quantum circuits and paves the way for future advancements in leveraging QML for medical imaging tasks.
2.3 Training quantum model on classical hardware
The training phase of our QML framework for medical image classification is executed on classical hardware as illustrated in Fig. 1. This approach circumvents the current limitations of quantum hardware in the NISQ era, such as limited qubits, high noise levels, and decoherence effects by performing the computationally expensive training process on classical GPUs even for large-scale datasets. Our QML workflow integrates classical optimization technique with noise-resilient, hardware-efficient quantum circuits tailored for medical imaging tasks. The training pipeline includes iterative optimization using gradient-based method and loss evaluation. The variational quantum circuit parameters $(\theta)$ are iteratively updated within a classical optimization loop to minimize a task-specific loss function. This iterative process continues until convergence is achieved or the maximum number of training epochs is reached, ensuring the parameters are optimized for the target classification task.
We conducted the training on noiseless simulators using the TorchQuantum library [41], which provides an ideal environment for parameter optimization without interference from hardware-induced noise. This approach significantly reduces computational overhead compared to direct hardware training and ensures consistent gradient computation. During each training iteration, the preprocessed input data is mapped to quantum states via encoding gates in the quantum circuit, which is then executed on the simulator. Expectation values obtained from circuit measurements are used to compute the loss function. Gradients of the loss function are calculated using the parameter-shift rule [30]. Afterwards, these gradients guide precise updates to the trainable parameters $(\theta)$. In our implementation, all the quantum circuits were trained for 200 epochs using the Adam optimizer with a learning rate of 0.01 and a batch size of 128, ensuring convergence to optimal parameter configurations. The MedMNIST contains datasets having different number of classes. During training, we measure one qubit and two qubits for two class and four class datasets, respectively and mean square loss function is used to calculate the loss. For the remaining multi-class datasets, we measure the number of qubits equals to the number of classes in the dataset and cross entropy loss function is used to calculate the loss. To ensure compatibility with MedMNIST datasets, we trained circuits with varying numbers of trainable parameters, such as 60, 80, 100, and 120 for each dataset of MedMNIST to account for differences in dataset complexity. The best-performing models are saved along with their parameters, losses, and accuracies for later inference on real quantum hardware. We also performed inference of these trained quantum models on simulator for all the considered MedMNIST datasets. All the quantum models are trained on a A100 Nvidia GPU of 43 GB RAM.
2.4 Inference on real IBM Cleveland hardware
The trained quantum models demonstrating the best classification performance on the simulator were selected for validation on the IBM Cleveland quantum hardware. During the inference stage of our QML framework, the trained quantum models with optimized parameters were utilized for evaluating the test samples. For each testing sample, the corresponding quantum circuit is dynamically constructed and transpiled at optimization level 3 to minimize circuit depth and gate count while ensuring compatibility with the hardware topology. Moreover, qubit connectivity was carefully configured to align with the hardware’s topology, minimizing the need for SWAP gates, which can introduce unnecessary noise and overhead. To further enhance the quantum model performance, we employed error suppression techniques, i.e. dynamical decoupling and gate twirling (Twir) during the sampling process on real IBM hardware. The DD is used to suppress the decoherence effects introduced due to the unwanted interaction between the qubits and the environment. Gate twirling applies random single-qubit gates around a target gate to average out coherent errors. The sampling on real hardware is performed using Qiskit SamplerV2 in Session execution mode. Moreover, we also integrated M3 error mitigation [42] in our QML workflow. M3 mitigation is a scalable quantum error correction approach that operates within the noisy bitstring subspace, correcting correlated and uncorrelated errors using a memory-efficient, matrix-free iterative solver. The integration of these error suppression and mitigation strategies enables quantum models to achieve optimal inference performance on real quantum hardware. Lastly, classical post-processing is applied on the mitigated bitstrings to obtain the predicted classification labels. To achieve this, the mitigated bitstrings are processed to compute qubit probabilities defined as:
$P(q,m)=\frac{N_{q,m}}{N_{\text{shots}}},\quad m\in\{0,1\}$ (1)
where $P(q,m)$ is the probability of measuring state $m$ on qubit $q$, $N_{q,m}$ is the number of times the $q$-th qubit is measured in state $\lvert m\rangle$ during the experiment. The measurement outcome $m=0$ corresponds to the state $\lvert 0\rangle$ and $m=1$ corresponds to the state $\lvert 1\rangle$. $N_{\text{shots}}$ represents the total number of measurement shots and we used $N_{\text{shots}}$ = 32000 for all the MedMNIST datasets. These probabilities provide the final expectation values for classification computed as:
$\langle Z_{q}\rangle=P(q,0)-P(q,1)$ (2)
where $\langle Z_{q}\rangle$ denotes the expectation value of the Pauli-Z operator for the $q$-th qubit, calculated as the difference between the probabilities of measuring 0 and 1. This expectation value vector is obtained for each test sample is processed to obtain the predicted class labels. The results from all test samples were aggregated to compute the performance metrices, providing insights into the performance of the quantum circuits on real quantum hardware. By conducting inference on real quantum device, we demonstrated the feasibility of deploying noise-resilient and device-aware quantum circuits for practical medical image classification tasks.
3 Experimental results
In this section, we conducted extensive experimental analysis to evaluate our QML workflow using both simulator and IBM Cleveland real hardware to benchmark MedMNIST datasets [1]. The MedMNIST contains 12 two-dimensional biomedical imaging datasets of different medical modalities and complexity levels. These 12 datasets includes 2-class as well as multi-class datasets including 4-class, 5-class, 7-class, 8-class, 9-class, and 11-class. For this benchmarking, we considered 8 datasets out of 12 by considering atleast one dataset from each category of different classes as shown in Table 1. The different medical modalities considered in these 8 datasets are illustrated in Fig. 2. We considered two widely used classification metrices i.e., accuracy (ACC) and area under the curve (AUC) to measure the performance of QML models. It is important to note that ACC provides a straightforward interpretation of model performance but it is highly sensitive to class imbalance, often leading to biased evaluations. On the contrary, AUC offers a more comprehensive evaluation of the classifier’s separability and is inherently robust to class imbalance. This makes AUC particularly suitable for datasets where class distributions are skewed. It is clear from the Table 1 that MedMNIST datasets are quite imbalance in nature. Therefore, we prioritized AUC over ACC to select model for real hardware execution. By focusing on AUC, we ensured that the chosen quantum models were able to effectively distinguish between classes across various thresholds, providing a fairer and more reliable assessment of performance.
Table 1: Data summary of MedMNIST-v2 datasets used in our QML workflow.
MedMNIST2D Data Modality Classes Training / Test Sample distribution
PneumoniaMNIST Chest X-Ray 2 4,708 / 624 Train: 1214, 3494; Test: 234, 390
BreastMNIST Breast Ultrasound 2 546 / 156 Train: 147, 399; Test: 42, 114
OCTMNIST Retinal OCT 4 97,477 / 1,000 Train: 33484, 10213, 7754, 46026; Test: 250, 250, 250, 250
RetinaMNIST Fundus Camera 5 1,080 / 400 Train: 486, 128, 206, 194, 66; Test: 174, 46, 92, 68, 20
DermaMNIST Dermatoscope 7 7,007 / 2,005
Train: 228, 359, 769, 80, 779, 4693, 99
Test: 66, 103, 220, 23, 223, 1341, 29
BloodMNIST Blood Cell Microscope 8 11,959 / 3,421
Train: 852, 2181, 1085, 2026, 849, 993, 2330, 1643
Test: 244, 624, 311, 579, 243, 284, 666, 470
PathMNIST Colon Pathology 9 89,996 / 7,180
Train: 9366, 9509, 10360, 10401, 8006, 12182, 7886, 9401, 12885
Test: 1338, 847, 339, 634, 1035, 592, 741, 421, 1233
OrganSMNIST Abdominal CT 11 13,932 / 8,827
Train: 1148, 630, 614, 721, 1132, 1119, 3464, 741, 803, 2004, 1556
Test: 811, 439, 445, 510, 704, 693, 2078, 397, 439, 1343, 968
Figure 2: MedMNIST datasets for 2D biomedical image classification.
The classical post-processing of M3-mitigated bitstrings in the inference stage provides expectation values vector for all test samples. These expectation values are used to calcuate the ACC and AUC metrices. For binary datasets (where we measure one qubit), the accuracy is calculated by measuring the proportion of predictions that deviate from the true labels by less than a threshold of 1. It evaluates how many predictions are close enough to the actual labels, providing an overall measure of correctness for the model’s outputs in binary classification tasks. The binary AUC is computed by assessing the separability of the two classes based on the predicted expectation values. We generates signed prediction labels to align with the true labels and calculates the AUC score using the AUC score function [43] from scikit-learn. For the multi-class datasets (where we are measuring more than 2 qubits), the model’s predicted class is determined by taking the index of the maximum value along each row of expectation values vector. This is compared against the true labels to compute the classification accuracy as:
ACC=Number of correct predictionsTotal number of $\text{ACC}=\frac{\text{Number of correct predictions}}{\text{Total number of samples}}$ (3)
To calculate the AUC metric for multi-class datasets, we first applied softmax activation function to convert the raw expectation values into class probabilities. This ensures that the outputs for each test sample sum to 1 and can be interpreted as probabilities for each class. Then, we use the AUC function [43] from scikit-learn to calculate the AUC.
Table 2: Performance analysis of QML framework on simulator based on classification accuracy (ACC) and area under the curve (AUC) metrices under different configurations of quantum models.
Dataset
Number of
qubits
Number of
embedding gates
Number of
trainable parameters
Noiseless
ACC
Noiseless
AUC
60 0.8478 0.8098
80 0.8462 0.8137
100 0.8462 0.8068
49 120 0.8526 0.8197
60 0.8253 0.785
80 0.8462 0.8128
100 0.8526 0.812
PneumoniaMNIST 4 64 120 0.8478 0.8132
60 0.769 0.652
80 0.801 0.684
100 0.814 0.674
49 120 0.808 0.693
60 0.776 0.681
80 0.769 0.722
100 0.775 0.658
BreastMNIST 4 64 120 0.794 0.694
60 0.433 0.657
80 0.41 0.656
100 0.404 0.648
49 120 0.415 0.657
60 0.388 0.628
80 0.409 0.631
100 0.423 0.651
4 64 120 0.413 0.648
60 0.387 0.62
80 0.379 0.623
100 0.443 0.659
49 120 0.441 0.652
60 0.393 0.621
80 0.36 0.593
100 0.398 0.631
OCTMNIST 8 64 120 0.379 0.628
60 0.4775 0.6268
80 0.48 0.6646
100 0.485 0.6434
49 120 0.5 0.6458
60 0.5025 0.6648
80 0.505 0.6474
100 0.5025 0.6586
RetinaMNIST 5 64 120 0.5175 0.6559
Table 3: Performance analysis of QML framework on simulator based on classification accuracy (ACC) and area under the curve (AUC) metrices under different configurations of quantum models.
Dataset
Number of
qubits
Number of
embedding gates
Number of
trainable parameters
Noiseless
ACC
Noiseless
AUC
60 0.6688 0.6873
80 0.6698 0.7144
100 0.6698 0.7013
49 120 0.6693 0.7059
60 0.6723 0.6938
80 0.6693 0.7094
100 0.6728 0.6959
DermaMNIST 7 64 120 0.6738 0.7343
60 0.415 0.774
80 0.371 0.781
100 0.462 0.811
49 120 0.432 0.782
60 0.44 0.792
80 0.434 0.796
100 0.421 0.791
BloodMNIST 8 64 120 0.482 0.821
60 0.3398 0.7542
80 0.3553 0.7492
100 0.3467 0.7703
49 120 0.3614 0.7796
60 0.3443 0.7574
80 0.3341 0.7708
100 0.3558 0.7682
PathMNIST 9 64 120 0.3723 0.784
60 0.3735 0.7492
80 0.3711 0.7585
100 0.369 0.7624
49 120 0.3544 0.7689
60 0.3353 0.729
80 0.3528 0.7709
100 0.3786 0.7599
OrganSMNIST 11 64 120 0.3681 0.7802
We present the noiseless results of the QML framework using simulator based on ACC and AUC metrices under different settings of input embedding features and quantum circuit trainable parameters for the considered MedMNIST datasets in Tables 2 and 3. The QML evaluation is performed by considering all the training and testing samples of MedMNIST datasets as given in Table 1. The bold values in the Tables 2 and 3 represent the best accuracy and AUC values. For PneumoniaMNIST dataset, our QML model with 4 qubits, 49 input embedding features, and 120 trainable parameters provides the highest classification accuracy of 85.26% and AUC of 81.97% as compared to the other settings as shown in Table 2. In the case of BreastMNIST dataset, we got best accuracy of 81.40% for the 4-qubit quantum model having 49 embedding features and 100 trainable parameters, while highest AUC of 72.20% is achieved for the model with 64 embedding features and 80 trainable parameters. Similarly, we got the highest accuracy of 44.30% and AUC of 65.90% for 8-qubit quantum model having 49 embedding features and 100 trainable parameters for OCTMNIST dataset. For RetinaMNIST dataset, we achieved best classification accuracy of 51.75% for the 5-qubit quantum model having 64 input embedding features and 120 trainable parameters. The highest AUC of 66.48% is obtained for the quantum model having 64 embedding features and 60 trainable parameters as depicted in Table 2. A 7-qubit quantum model with 64 input embedding features and 120 trainable parameters provides highest accuracy of 67.38% and AUC of 73.43% for the 7-class DermaMNIST dataset as shown in Table 3. For a 8-class BloodMNIST dataset, we achieved the highest accuracy of 48.20% and AUC of 82.10% using the QML model with 64 embedding features and 120 trainable parameters. Similarly, we got the best accuracy of 37.23% and AUC of 78.40% for a 9-class PathMNIST dataset using the QML model having 64 embedding features and 120 trainable parameters. Lastly, the QML model with 64 features and 100 trainable parameters provides best classification accuracy of 37.86% for 11-class OrganSMNIST dataset, while best AUC of 78.02% is achieved for the QML model having 64 embedding features and 120 trainable parameters. It is clear from these results that the quantum models without any classical neural networks part and with very limited input features showing promising capability to classify medical imaging datasets. Also, it is observed from the Tables 2 and 3 that there is a significant difference between the accuracy and AUC values, for instance, we got an accuracy of 48.20% and AUC of 82.10% for BloodMNIST dataset. Similar trend can be seen for the PathMNIST and OrganSMNIST datasets. Note that we considered AUC as our primary performance evaluation metric because it is less sensitive to the class imbalance of MedMNIST datasets as compared to the accuracy metric. Moreover, it can be noted that for most of the multi-class datasets, we got best results corresponding to the QML models having more number of embedding features and trainable parameters in the quantum circuit.
Table 4: Ablation of different error suppression/mitigation techniques using OCTMNIST dataset on real IBM Cleveland hardware.
IBM Cleveland (ACC) IBM Cleveland (AUC)
No error suppression/mitigation 0.286 0.545
Error suppression (DD+Twirling) 0.33 0.591
M3 error mitigation 0.285 0.533
DD+Twirling+M3 0.397 0.622
We also conducted an ablation study on IBM Cleveland real hardware to understand the impact of error suppression and error mitigation techniques on the performance of QML models for medical imaging classification as shown in Table 4. We performed this ablation on the OCTMNIST dataset because of the significant effect of hardware noise on the performance of QML model. For OCTMNIST dataset, our QML workflow provides the noiseless accuracy of 44.30% and AUC of 65.90% as shown in Table 2. When we run this quantum model on real IBM hardware without any error suppression and mitigation techniques, we got an accuracy of 28.60% and AUC of 54.50%. Therefore, we observed the reduction of 15.70% and 11.40% in accuracy and AUC, respectively due to the noise of IBM Cleveland hardware. In this ablation, since we are using sampler in the inference stage as shown in Fig. 1, we considered two error suppression techniques i.e., dynamical decoupling and gate twirling and one error mitigation technique i.e., M3 mitigation. It is observed from the Table 4 that when we use both error suppression techniques and M3 mitigation together (DD+Twirling+M3), we got best accuracy of 39.70% and AUC of 62.20%, thereby providing an improvements of 11.10% and 7.7% in accuracy and AUC respectively as compared to the case when no suppression and mitigation is used. Based on this ablation study on OCTMNIST dataset, we used this combination (DD+Twirling+M3) for all the considered MedMNIST datasets.
Table 5: Performance analysis of QML framework on real IBM Cleveland hardware based on classification accuracy (ACC) and area under the curve (AUC) metrices for different MedMNIST datasets.
Dataset Qubits Embed gates Tunable params Circuit gates Circuit depth Noiseless IBM Cleveland
No supp./miti. DD+Twir+M3
ACC AUC ACC AUC ACC AUC
PneumoniaMNIST 4 49 120 184 51 0.8526 0.8197 0.8381 0.8038 0.8542 0.8218
BreastMNIST 4 64 80 187 59 0.769 0.722 0.782 0.738 0.769 0.714
OCTMNIST 8 49 100 208 47 0.443 0.659 0.286 0.545 0.397 0.622
RetinaMNIST 5 64 60 140 29 0.5025 0.6648 0.5025 0.6638 0.505 0.6642
DermaMNIST 7 64 120 212 33 0.6738 0.7343 0.671 0.6825 0.677 0.6982
BloodMNIST 8 64 120 211 27 0.482 0.821 0.422 0.824 0.424 0.826
PathMNIST 9 64 120 218 27 0.3723 0.784 0.35 0.7544 0.418 0.7795
OrganSMNIST 11 64 120 213 23 0.3681 0.7802 0.227 0.7322 0.34 0.7678
We selected the best performing QML models from Tables 2 and 3 based on the AUC metric for inference on real IBM Cleveland hardware for all the considered MedMNIST datasets. Due to the hardware limitations, we considered 1000 testing samples to evaluate the QML models for the datasets having large number of testing samples. We presented noiseless and IBM hardware results (with and without error suppression and mitigation) together in Table 5 for better understanding along with the quantum models specifications such as number of qubits, number of embedded features/gates, trainable parameters, quantum circut gates, and circuit depth. It is observed from the Table 5 that even with the small number of input embedding features, our QML models consistently provide promising results in terms of classification accuracy and AUC metrices. It is worth noting that the noiseless results and hardware results (without error suppression and mitigation) are quite comparable for most of the MedMNIST datasets except OCTMNIST and OrganSMNIST dataset. This is because we generated device-aware quantum circuits with optimal circuit gates, depth, and robust against hardware noises. For OCTMNIST dataset, the classification accuracy and AUC on the hardware (without error suppression and mitigation) reduce by 15.70% and 11.40% respectively as compared to the noiseless results. Similarly, accuracy and AUC decrease by 14.11% and 4.8% respectively for OrganSMNIST dataset as compared to the noiseless results. This indicates that there is scope for further improvement in the design of device-aware quantum circuits. To mitigate the effect of hardware noises, we run the hardware experiments for all the considered MedMNIST datasets using error suppression and error mitigation combination (DD+Twirling+M3) based on the ablation study conducted in Table 4. We noted improvements in classification accuracy and AUC results using DD+Twirling+M3 combination as compared to the hardware results without error suppression and mitigation for all the datasets except BreastMNIST dataset. For instance, the classification accuracy and AUC values are significantly improved by 11.10% and 7.7% respectively as compared to the hardware results without error suppression and mitigation for OCTMNIST dataset. For BreastMNIST dataset, the accuracy and AUC values decreased by 1.3% and 2.4% respectively as compared to the hardware results without error suppression and mitigation.
It is important to mention that all classical ML models leverage the full spatial dimensions of input images, whereas our QML model operates on a significantly reduced feature set, dictated by quantum hardware constraints. Despite this, our QML model achieves promising results on real IBM cleveland hardware in terms of classification accuracy and AUC values. For the 2-class PneumoniaMNIST dataset, we got very good classification accuracy and AUC of 85.42% and 82.18%, respectively. Similarly, we achieved an accuracy and AUC of 76.90% and 71.40%, respectively in the case of 2-class BreastMNIST dataset. Even for 7-class DermaMNIST dataset, our QML model obtained very good results with an accuracy and AUC values of 67.70% and 69.82%, respectively. Our QML model is able to achieve the AUC of more than 70% for all considered MedMNIST datasets except OCTMNIST, RetinaMNIST, and DermaMNIST datasets. It is worth mentioning that for 2-class PneumoniaMNIST and 8-class BloodMNIST datasets, we got the AUC of more than 80% on real quantum hardware. To the best of our knowledge, this benchmarking study is the first to provide a comprehensive evaluation of QML (without using any classical neural networks) on real IBM quantum hardware considering large-scale MedMNIST datasets. Our study provides a critical evaluation of QML for realistic applications, showcasing its capability even in the evolving phase of quantum computing. By systematically benchmarking QML performance on real IBM Cleveland hardware across diverse MedMNIST datasets, this work paves the way for future advancements in QML and highlights its potential to complement classical ML in the years to come.
4 Conclusion
This work demonstrates the potential of QML in the context of medical imaging by benchmarking the MedMNIST datasets on real quantum hardware. We effectively trained and validated noise-resilient quantum circuits and leveraging advanced error suppression and mitigation techniques to tackle the hardware noise. The experimental results highlight the feasibility of applying QML to complex real-world datasets such as MedMNIST. In the current NISQ era, we are able to achieve promising results in terms of classification accuracy and AUC values for binary as well as multi-class datasets with very limited number of input embedding features. This study serves as an important benchmark, demonstrating the readiness of quantum computing for practical applications in healthcare. It also underscores the need for further research to enhance scalability, mitigate hardware limitations, and explore novel quantum algorithms tailored to medical data. As quantum technology continues to evolve, its integration into medical imaging workflows has the potential to transform healthcare analytics, paving the way for more advanced and efficient diagnostic tools.
Acknowledgments
The authors gratefully acknowledge financial support from the NIH (GM130641). We also extend our sincere gratitude to Danil Kaliakin (Research Associate, Cleveland Clinic) and Abdullah Ash Saki (Researcher at IBM Quantum) for their valuable insights and support throughout this project. We are also grateful to the authors of Élivágar and Torchquantum for their great work and open-sourcing the code.